A study to determine the relationship between the height of a son and the height of his mother and father used a random sample of size n=15 with a multiple regression equation was =22.97+.30x1+.41x2where y is the height of the son, x1is the height of the mother and x2is the height of the father.  The standard error coefficient for B1is .06107 and for B2is .0544.

a.  Predict the height of the son, if the mother’s height is 70 inches and the father’s height is 72 inches.

      b.   Find the 90% confidence interval for the coefficient B1.
Answer #1

a)predicted height =22.97+.30*70+.41*72 =73.49

b)

sample size n= 15
number of independent variables p= 2
degree of freedom =n-p-1= 12
estimated slope b= 0.3
standard error of slope=sb=sqrt(SSE/(n-p-1))/sqrt(SSx)= 0.061070
for 90 % confidence and 12degree of freedom critical t= 1.7820
90% confidence interval =b1 -/+ t*standard error= (0.1912,0.4088)